A steel casting initially at 0 °C is placed in a furnace at 800 C°. If it increases in temperature at the
rate of 5.00% per minute, find its temperature after 20.0 min.

Found 2 solutions by ikleyn, Alan3354:
Answer by ikleyn(52943)

.

To find the answer, multiply the rate by the time.


                                ////////////


The post by Alan, which he wrote in response to my post, is   T O T A L L Y   A B S U R D I S T,
and I should explain  WHY.

First of all, there is a mistake (a typo) in the original post.

To be reasonable and accurate, it must say "the temperature increase rate is 5° C per minute", and not "5.00% per minute".

It is SO OBVIOUS, that I did not explain or fix it in my previous post . . .


After that, it becomes clear that the temperature increase is the product of the rate of the temperature increase by the time,
as it is written in my previous post.


So, the temperature increase is 5* 20 = 100 degrees of Celsius, and the final temperature is 0° C + 5*20° C o° + 100°C = 100° C.


Again, it is at the level of knowledge, when a student learn the conception of temperature for the first time in his (or her) life.


So, using Kelvin's temperature and using exponential function for the temperature is ABSOLUTELY IRRELEVANT.


Alan, honestly, I did not expect to see such a "solution" from you.


            The temperature of the solid may rise exponentially  ONLY  IN  ONE  CASE.

                    It is the case, when there is a chain reaction inside of it,
                which creates an interior source of heat,  like a radioactive decay,
                        which in some extreme cases is a nuclear explosion.



Answer by Alan3354(69443)

A steel casting initially at 0 °C is placed in a furnace at 800 C°. If it increases in temperature at the
rate of 5.00% per minute, find its temperature after 20.0 min.
======================
Degs C won't work. 5% of zero is zero.
Use Kelvin.
0 C =~ 273K
---
After 20 mins, T = 273*1.05^20 = 724K
= 451C